# Context-Alignment: Activating and Enhancing LLM Capabilities in Time Series

### v1

@@ 4 Experiments
The proposed Context-Alignment demonstrates robust performance across a variety of tasks, detailed in Sec. 4.2 (long-term forecasting), Sec. 4.3 (short-term forecasting), Sec. 4.4 (few-shot forecasting), Sec. 4.5 (zero-shot forecasting), and Sec. 4.6 (classification). Most experiments leverage DECA. Specifically, VCA validates its efficacy using logic guidance and structural division alone in section 4.1. For classification tasks with multiple classes, where GPT-2’s length constraints are limiting, DECA is reserved for binary classes, and VCA is applied to multi-class datasets. Context-Alignment significantly boosts training efficiency and cost-effectiveness, especially in few-shot and zero-shot forecasting, by establishing robust a priori structural understanding and logical relationships. Due to space limitations, comprehensive results are detailed in Appendix C.
Baselines.  Building on the groundwork of Zhou et al. (2023) and Jin et al. (2024), and mindful of page constraints, we selected a representative array of high-performing baselines for extensive evaluation. These encompass Transformer-based models such as iTransformer Liu et al. (2023a), FEDformer Zhou et al. (2022), Non-stationary Transformer Liu et al. (2022), ETSformer Woo et al. (2022), PatchTST Nie et al. (2022), alongside notable non-Transformer methods like TimesNet Wu et al. (2022), and DLinear Zeng et al. (2023). We also included advanced LLM-based models—GPT4TS Zhou et al. (2023), Time-LLM Jin et al. (2024), and S2IP-LLM Pan et al. (2024)—all utilizing GPT-2 as the standard LLM backbone to ensure model consistency. To ensure a fair comparison, we adhere to the experimental framework outlined in  Zhou et al. (2023) and  Wu et al. (2022). Detailed evaluations are expanded upon in subsequent sections.
@@ 4.1 Demo: Vanilla Context-Alignment
VCA achieves logical and structural alignment through DSCA-GNNs. As demonstrated on the ETT dataset (Table 1), VCA substantially outperforms both variant without DSCA-GNNs and other baselines. A performance decline is observed compared to DECA, which further augments LLMs’ TS processing capabilities using demonstration examples prompt. VCA without DSCA-GNNs, despite incorporating language prompts, lacks context-alignment GNNs, resulting in more verbose and semantically confused inputs to the LLM, thus yielding the worst outcomes.
@@ 4.2 Long-term Forecasting
Setups.  For long-term forecasting task, we validate the efficacy of DECA across eight prevalent datasets Wu et al. (2022): ETTh1, ETTh2, ETTm1, ETTm2, Weather, Electricity, Traffic, and ILI. Consistent with GPT4TS Zhou et al. (2023), Time-LLM Jin et al. (2024), and S2IP-LLM Pan et al. (2024), we utilize an input TS length of 512 except for the ILI dataset. Performance is evaluated over four prediction horizons: $\{24,36,48,60\}$ for ILI and $\{96,192,336,720\}$ for the other datasets, using Mean Squared Error (MSE) and Mean Absolute Error (MAE) as metrics.
Results.  As shown in Table 2, DECA surpasses all baseline methods in most scenarios. Specifically, it reduces average MSE of 3.1% over the suboptimal method PatchTST, and outperforms other LLM-based methods—S2IP-LLM, Time-LLM, and GPT4TS by 7.3%, 12.2%, and 16.6%. This consistent superiority across diverse datasets highlights the critical role of logical and structural alignment. Furthermore, demonstration examples prompt boost LLMs’ contextual understanding of TS data.
@@
width = 1colsep = 2.2pt, rowsep = 1.4pt, colspec = Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c], cells = c, cell11,3,5,7,9,11,13,15,17,19,21 = c=2, cell1,2,3,4,5,6,7,8,9,101 = c=2, hline1 = 1-240.15em, hline2,3 = 1-240.05em, hline11 = 1-240.15em, vline2,3,5,7,9,11,13,15,17,19,21,23 = 4-9, vline2,3,5,7,9,11,13,15,17,19,21,23 = 3abovepos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 10belowpos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 1,2abovepos = -1.2, belowpos = -1.2, Methods & DECA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST TimesNet FEDformer Stationary ETSformer
@@ 4.3 Short-term Forecasting
width = 1colspec = Q[0.2em,c]Q[2em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c], colsep = 0.1pt, rowsep = 1.4pt, cell11 = c=20.09, cell21 = r=3, hline1 = 1-140.15em, hline2 = 1-140.05em, hline5 = 1-140.15em, vline3,4,5,6,7,8,9,10,11,12,13,14 = 3, vline3,4,5,6,7,8,9,10,11,12,13,14 = 2abovepos = -1.2, vline3,4,5,6,7,8,9,10,11,12,13,14 = 4belowpos = -1.2, vline3,4,5,6,7,8,9,10,11,12,13,14 = 1abovepos = -1.2, belowpos = -1.2, Methods & DECA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST N-HiTS N-BEATS TimesNet FEDformer Stationary
@@
Setups.  We conduct short-term forecasting experiments on the M4 dataset Makridakis et al. (2018), which includes market data across various frequencies, with prediction horizons ranging from 6 to 48. We incorporate N-HiTS Challu et al. (2023) and N-BEATS Oreshkin et al. (2019) as additional baselines. Performance is quantified using symmetric mean absolute percentage error (SMAPE), mean absolute scaled error (MASE), and the overall weighted average (OWA) as evaluation metrics.
Results.  Results in Table 3 indicate that DECA exhibits competitive performance compared to SOTA methods. DECA maintains robustness in both long-term and short-term forecasting, attributable to the effectiveness of structural and logical alignment across varying sequence lengths.
@@ 4.4 Few-shot Forecasting
width = 1colspec = Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c], colsep = 2.2pt, rowsep = 1.4pt, cells = c, cell11,3,5,7,9,11,13,15,17,19,21,23 = c=2, cell2,3,4,5,6,71 = c=2, hline1,8 = 1-240.15em, hline2,3,7 = 1-240.05em, vline2,3,5,7,9,11,13,15,17,19,21,23 = 4-5, vline2,3,5,7,9,11,13,15,17,19,21,23 = 3abovepos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 6belowpos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 1,2,7abovepos = -1.2, belowpos = -1.2, Methods & DECA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST TimesNet FEDformer Stationary ETSformer
@@
Setups.  The expressive potential of LLMs often results in robust performance in few-shot scenarios Brown (2020); Liu et al. (2023b). While current LLM-based methods outperform earlier models like DLinear, PatchTST, and TimesNet, they do not fully exploit LLMs’ deep TS comprehension, thus not maximizing their few-shot capabilities. To explore whether Context-Alignment can boost the few-shot efficacy of LLMs, we perform experiments on the ETT datasets, following the protocol established by Jin et al. (2024). We evaluate the performance of DECA using only 5% of training data, with results from 10% data detailed in Appendix C.3.
Results.  As shown in Table 4, our method consistently outperforms all baselines. It achieves a 6.7% reduction in average MSE compared to the leading LLM-based model, S2IP-LLM. We observe further improvements of 9.4% and 10.8% against Time-LLM and GPT4TS. Moreover, DECA shows a 15.8% performance gain over the SOTA transformer model PatchTST. We attribute these to the integration of prior knowledge in structural division and logic guidance by Context-Alignment. The strong generalizability of these priors, even with limited training data, effectively activates LLMs’ latent few-shot capabilities in TS, further boosted by a few demonstration examples.
@@ 4.5 Zero-shot Forecasting
Setups.  Despite the inherent zero-shot capabilities of LLMs Kojima et al. (2022), LLM-based methods struggle to fundamentally understand the structure of TS data and its logical associations with prompts, leading to underperformance compared to Transformer-based methods like PatchTST. Using the setup from Time-LLM Jin et al. (2024), we also evaluate the cross-domain efficacy of DECA on ETT dataset. Specifically, the model is trained on Dataset A and then tested on Dataset B without utilizing any training data from Dataset B.
@@
Methods & DECA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST TimesNet
@@
Figure 2: For the classification task, the results show the average accuracy across 10 subsets from the UEA dataset. The complete results are displayed in Appendix C.5.
Results.  Table 5 shows that DECA substantially outperforms the most competitive baselines. Compared to the second-best model, PatchTST, DECA exhibits a 13.3% improvement in performance. Against LLM-based models like S2IP-LLM, Time-LLM, and GPT4TS, DECA achieves performance gains of 18.3%, 17.7%, 24.3%. Experiments in both few-shot and zero-shot settings highlight DECA’s exceptional performance under data-scarce conditions. In these scenarios, Context-Alignment paradigm provides a robust contextual prior, enabling accurate logical and structural understanding that enhances the potential of LLMs for cross-domain TS processing.
@@ 4.6 Time Series Classification
Setups.  To assess the model’s capability in learning advanced representations, we conduct comparative experiments on series classification using the setup outlined by Zhou et al. (2023). We select 10 multivariate UEA classification datasets Bagnall et al. (2018) from domains including ECG, audio, gesture, spectral recognition, etc. For binary class datasets like FaceDetection, we apply DECA framework, adding a demonstration example prompt for each category at the beginning of the prediction sequence. For datasets with multi-class, such as Handwriting with 26 classes, we employ VCA due to GPT-2’s input length constraints.
Results.  Fig. 2 presents the average accuracy of various methods on all UEA classification datasets. DECA achieves an accuracy of 76.4%, surpassing all baseline methods and recording a 2.4% increase over the next best model, Zhou et al. (2023). This performance suggests the effectiveness of DECA or VCA extends beyond predictive tasks. Additionally, Context-Alignment shows generality, indicating its potential applicability across various contexts in the field.
@@ 4.7 Ablation Study
As shown in Table 6, we conduct an ablation study on the framework design and analyze the effectiveness of DECA. DECA represents the optimal results, highlighted in bold in Table 6. D.4 is optimal for long-term forecasting, whereas D.3 is best for classification tasks.
Validity of Dual-Scale Context-Alignment GNNs.  To evaluate DSCA-GNNs, we conduct comparative experiments from two perspectives. In A.1, without DSCA-GNNs, we rely solely on demonstration examples as prompts, resulting in higher average MSE compared to DECA. In A.2, random initialization of adjacency matrix leads to further performance decline compared to A.1, underscoring that incorrect logical information can impair model performance. A comprehensive comparison across A.1, A.2, and DECA confirms that GNNs guided by appropriate logical frameworks effectively leverages the inherent capabilities of LLMs in TS tasks.
@@
B.1 w/o Coarse-grained GNN 0.401 0.353 69.6 78.5
C.1 DECA(2) 0.402 0.357 69.1 78.0
C.2 DECA(4) 0.396 0.345 70.4 79.5
@@
DECA(6) 0.399 0.347 70.1 79.0
C.4 DECA(8) 0.418 0.362 67.7 77.0
C.5 DECA(10) 0.439 0.383 66.4 75.6
C.6 DECA(12) 0.455 0.396 63.2 72.6
@@
Validity of Coarse-Grained GNN.  In B.1, omitting this module impairs the model’s ability to understand the overall structure and the macro-level logical relationships, resulting in the reduced performance compared to DECA. This decline is due to reliance solely on the fine-grained GNN for token-scale context-alignment, which still leaves the inputs verbose and lacking structure.
The Number of LLM Layers.  We conduct ablation studies to assess the impact of varying numbers of GPT-2 layers. C.* indicates that models with 4 and 6 layers perform optimally, while adding more layers causes overfitting Zhou et al. (2023). For computational efficiency, we select 4 layers (C.2).
Insertion Position of Dual-Scale Context-Alignment GNNs.  DSCA-GNNs can integrate seamlessly into various layers of LLMs. Firstly, when priors for structural and logical alignment are present at the input, deeper layers of LLMs can more fully leverage this prior knowledge, hence, we default to employing this module at the model’s input, as shown in configuration D.1. Secondly, D.2, D.3, and D.4 involve repeating integration of this module at mid and output stages of the LLMs encoder, enhancing prior guidance and yielding performance improvements. Ablation studies for long-term forecasting and classification tasks reveal that optimal performance results from different insertion points, attributed to domain differences in token features among various LLMs layers Ethayarajh (2019) and varying demands of different tasks; Thirdly, configuration D.5, which omits GNNs module at model’s input, leads to decreased performance, confirming that incorporating context-alignment at initial stage enhances the overall utilization of pretrained LLMs.

### v2

@@ 4 Experiments
The proposed Context-Alignment demonstrates robust performance across a variety of tasks, detailed in Sec. 4.2 (long-term forecasting), Sec. 4.3 (short-term forecasting), Sec. 4.4 (few-shot forecasting), Sec. 4.5 (zero-shot forecasting), and Sec. 4.6 (classification). Most experiments leverage FSCA. Specifically, VCA validates its efficacy using logic guidance and structural division alone in section 4.1. For classification tasks with multiple classes, where GPT-2’s length constraints are limiting, FSCA is reserved for binary classes, and VCA is applied to multi-class datasets. Context-Alignment significantly boosts training efficiency and cost-effectiveness, especially in few-shot and zero-shot forecasting, by establishing robust a priori structural understanding and logical relationships. Full results and dataset description are in Appendix C and A.2, respectively.
Baselines.  Building on the groundwork of Zhou et al. (2023) and Jin et al. (2024), and mindful of page constraints, we selected a representative array of high-performing baselines for extensive evaluation. These encompass Transformer-based models such as iTransformer (Liu et al., 2023a), FEDformer (Zhou et al., 2022), Non-stationary Transformer (Liu et al., 2022), ETSformer (Woo et al., 2022), PatchTST (Nie et al., 2022), alongside notable non-Transformer methods like TimesNet (Wu et al., 2022), and DLinear (Zeng et al., 2023). We also included advanced LLM-based models—GPT4TS (Zhou et al., 2023), Time-LLM (Jin et al., 2024), and S2IP-LLM (Pan et al., 2024)—all utilizing GPT-2 as the standard LLM backbone to ensure model consistency. To ensure a fair comparison, we adhere to the experimental framework outlined in  Zhou et al. (2023) and  Wu et al. (2022). Detailed evaluations are expanded upon in subsequent sections.
@@ 4.1 Demo: Vanilla Context-Alignment
VCA achieves logical and structural alignment through DSCA-GNNs. As demonstrated on the ETT dataset (Table 1), VCA substantially outperforms both variants without DSCA-GNNs and other baselines. A performance decline is observed compared to FSCA, which further augments LLMs’ TS processing capabilities using demonstration examples prompt. VCA without DSCA-GNNs, despite incorporating task description prompts, lacks context-alignment GNNs, resulting in more verbose and semantically confused inputs for LLM, thus yielding the worst outcomes.
@@ 4.2 Long-term Forecasting
Setups.  For long-term forecasting tasks, we validate the efficacy of FSCA across eight prevalent datasets (Wu et al., 2022): ETTh1, ETTh2, ETTm1, ETTm2, Weather, Electricity, Traffic, and ILI. Consistent with GPT4TS (Zhou et al., 2023), Time-LLM (Jin et al., 2024), and S2IP-LLM (Pan et al., 2024), we utilize an input TS length of 512 except for the ILI dataset. Performance is evaluated over four prediction horizons: $\{24,36,48,60\}$ for ILI and $\{96,192,336,720\}$ for the other datasets, using Mean Squared Error (MSE) and Mean Absolute Error (MAE) as metrics.
Results.  As shown in Table 2, FSCA surpasses all baseline methods in most scenarios. Specifically, it reduces average MSE of 3.1% over the suboptimal method PatchTST, and outperforms other LLM-based methods—S2IP-LLM, Time-LLM, and GPT4TS by 7.3%, 12.2%, and 16.6%. This consistent superiority across diverse datasets highlights the critical role of logical and structural alignment. Furthermore, demonstration examples prompt boost LLMs’ contextual understanding of TS data.
@@
width = 1colsep = 2.2pt, rowsep = 1.4pt, colspec = Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c], cells = c, cell11,3,5,7,9,11,13,15,17,19,21 = c=2, cell1,2,3,4,5,6,7,8,9,101 = c=2, hline1 = 1-240.15em, hline2,3 = 1-240.05em, hline11 = 1-240.15em, vline2,3,5,7,9,11,13,15,17,19,21,23 = 4-9, vline2,3,5,7,9,11,13,15,17,19,21,23 = 3abovepos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 10belowpos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 1,2abovepos = -1.2, belowpos = -1.2, Methods & FSCA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST TimesNet FEDformer Stationary ETSformer
@@ 4.3 Short-term Forecasting
width = 1colspec = Q[0.2em,c]Q[2em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c], colsep = 0.1pt, rowsep = 1.4pt, cell11 = c=20.09, cell21 = r=3, hline1 = 1-140.15em, hline2 = 1-140.05em, hline5 = 1-140.15em, vline3,4,5,6,7,8,9,10,11,12,13,14 = 3, vline3,4,5,6,7,8,9,10,11,12,13,14 = 2abovepos = -1.2, vline3,4,5,6,7,8,9,10,11,12,13,14 = 4belowpos = -1.2, vline3,4,5,6,7,8,9,10,11,12,13,14 = 1abovepos = -1.2, belowpos = -1.2, Methods & FSCA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST N-HiTS N-BEATS TimesNet FEDformer Stationary
@@
Setups.  We conduct short-term forecasting experiments on the M4 dataset (Makridakis et al., 2018), which includes market data across various frequencies, with prediction horizons ranging from 6 to 48. We incorporate N-HiTS (Challu et al., 2023) and N-BEATS (Oreshkin et al., 2019) as additional baselines. Performance is quantified using symmetric mean absolute percentage error (SMAPE), mean absolute scaled error (MASE), and the overall weighted average (OWA) as evaluation metrics.
Results.  Results in Table 3 indicate that FSCA exhibits competitive performance compared to SOTA methods. FSCA maintains robustness in both long-term and short-term forecasting, attributable to the effectiveness of structural and logical alignment across varying sequence lengths.
@@ 4.4 Few-shot Forecasting
width = 1colspec = Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c], colsep = 2.2pt, rowsep = 1.4pt, cells = c, cell11,3,5,7,9,11,13,15,17,19,21,23 = c=2, cell2,3,4,5,6,71 = c=2, hline1,8 = 1-240.15em, hline2,3,7 = 1-240.05em, vline2,3,5,7,9,11,13,15,17,19,21,23 = 4-5, vline2,3,5,7,9,11,13,15,17,19,21,23 = 3abovepos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 6belowpos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 1,2,7abovepos = -1.2, belowpos = -1.2, Methods & FSCA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST TimesNet FEDformer Stationary ETSformer
@@
Setups.  The expressive potential of LLMs often results in a robust performance in few-shot scenarios (Brown, 2020; Liu et al., 2023b). While current LLM-based methods outperform earlier models like DLinear, PatchTST, and TimesNet, they do not fully exploit LLMs’ deep TS comprehension, thus not maximizing their few-shot capabilities. To explore whether Context-Alignment can boost the few-shot efficacy of LLMs, we perform experiments on the ETT datasets, following the protocol established by Jin et al. (2024). We evaluate the performance of FSCA using only 5% of training data, with results from 10% data detailed in Appendix C.3.
Results.  As shown in Table 4, our method consistently outperforms all baselines. It achieves a 6.7% reduction in average MSE compared to the leading LLM-based model, S2IP-LLM. We observe further improvements of 9.4% and 10.8% against Time-LLM and GPT4TS. Moreover, FSCA shows a 15.8% performance gain over the SOTA transformer model PatchTST. We attribute these to the integration of prior knowledge in structural division and logic guidance by Context-Alignment. The strong generalizability of these priors, even with limited training data, effectively activates LLMs’ latent few-shot capabilities in TS, further boosted by a few demonstration examples.
@@ 4.5 Zero-shot Forecasting
Setups.  Despite the inherent zero-shot capabilities of LLMs (Kojima et al., 2022), LLM-based methods struggle to fundamentally understand the structure of TS data and its logical associations with prompts, leading to underperformance compared to Transformer-based methods like PatchTST. Using the setup from Time-LLM (Jin et al., 2024), we also evaluate the cross-domain efficacy of FSCA on ETT dataset. Specifically, the model is trained on Dataset A and then tested on Dataset B without utilizing any training data from Dataset B.
@@
Methods & FSCA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST TimesNet
@@
Results.  Table 5 shows that FSCA substantially outperforms the most competitive baselines. Compared to the second-best model, PatchTST, FSCA exhibits a 13.3% improvement in performance. Against LLM-based models like S2IP-LLM, Time-LLM, and GPT4TS, FSCA achieves performance gains of 18.3%, 17.7%, 24.3%. Experiments in both few-shot and zero-shot settings highlight FSCA’s exceptional performance under data-scarce conditions. In these scenarios, Context-Alignment paradigm provides a robust contextual prior, enabling accurate logical and structural understanding that enhances the potential of LLMs for cross-domain TS processing.
@@ 4.6 Time Series Classification
Figure 2: For the classification task, the results show average accuracy across 10 subsets from the UEA dataset. Detailed results are in Appendix F.1.
Setups.  To assess the model’s capability in learning advanced representations, we conduct comparative experiments on series classification using the setup outlined by Zhou et al. (2023). We select 10 multivariate UEA classification datasets (Bagnall et al., 2018) from domains including ECG, audio, gesture, spectral recognition, etc. For binary class datasets like FaceDetection, we apply FSCA framework, adding a demonstration example prompt for each category at the beginning of the prediction sequence. For datasets with multiple classes, such as Handwriting with 26 classes, we employ VCA due to GPT-2’s input length constraints.
Results.  Fig. 2 presents the average accuracy of various methods on all UEA classification datasets. FSCA achieves an accuracy of 76.4%, surpassing all baseline methods and recording a 2.4% increase over the next best model, Zhou et al. (2023). This performance suggests the effectiveness of FSCA or VCA extends beyond predictive tasks. Additionally, Context-Alignment shows generality, indicating its potential applicability across various contexts in the field.
@@ 4.7 Ablation Study
As shown in Table 6, we conduct an ablation study on the framework design and analyze the effectiveness of FSCA. FSCA represents the optimal results, highlighted in bold in Table 6. D.4 is optimal for long-term forecasting, whereas D.3 is best for classification tasks.
Validity of Dual-Scale Context-Alignment GNNs.  To evaluate DSCA-GNNs, we conduct comparative experiments from two perspectives. In A.1, without Dual-Scale GNNs, we rely solely on demonstration examples as prompts, resulting in higher average MSE compared to FSCA. In A.2, random initialization of adjacency matrix leads to further performance decline compared to A.1, underscoring that incorrect logical information can impair model performance. A comprehensive comparison across A.1, A.2, and FSCA confirms that GNNs guided by appropriate logical frameworks effectively leverage the inherent capabilities of LLMs in TS tasks.
Validity of Coarse-Grained Branch.  In B.1, omitting this module impairs the model’s ability to understand the overall structure and the macro-level logical relationships, resulting in reduced performance compared to FSCA. This decline is due to reliance solely on the fine-grained branch for token-scale context alignment, which still leaves the inputs verbose and lacking structure.
The Number of LLM Layers.  We conduct ablation studies to assess the impact of varying numbers of GPT-2 layers. C.* indicates that models with 4 and 6 layers perform optimally while adding more layers causes overfitting (Zhou et al., 2023). For computational efficiency, we select 4 layers (C.2).
@@
B.1 w/o Coarse-grained Branch 0.401 0.353 69.6 78.5
C.1 FSCA(2) 0.402 0.357 69.1 78.0
C.2 FSCA(4) 0.396 0.345 70.4 79.5
@@
FSCA(6) 0.399 0.347 70.1 79.0
C.4 FSCA(8) 0.418 0.362 67.7 77.0
C.5 FSCA(10) 0.439 0.383 66.4 75.6
C.6 FSCA(12) 0.455 0.396 63.2 72.6
@@
Insertion Position of Dual-Scale Context-Alignment GNNs.  DSCA-GNNs can seamlessly integrate into various layers of LLMs. Firstly, when priors for structural and logical alignment are present at the input, deeper layers of LLMs can more fully leverage this prior knowledge, hence, we default to employing this module at the model’s input, as shown in configuration D.1. Secondly, D.2, D.3, and D.4 involve repeating integration of this module at the mid and output stages of the LLMs encoder, enhancing prior guidance and yielding performance improvements. Ablation studies for long-term forecasting and classification tasks reveal that optimal performance results from different insertion points, attributed to domain differences in token features among various LLMs layers (Ethayarajh, 2019) and varying demands of different tasks; Thirdly, configuration D.5, which omits GNNs module at model’s input, leads to decreased performance, confirming that incorporating context-alignment at initial stage enhances the overall utilization of pre-trained LLMs.